OpenStudy (rayman5766):

Which expressions, when evaluated with z = 8, are negative in value? Choose all answers that are correct. A. –2z – 14 B. –25 + 4z C. –2z + 15 D. 30 – 3z

OpenStudy (rayman5766):

keep it up and I will become your fan

OpenStudy (haichi):

Medal plz

OpenStudy (haichi):

kk

OpenStudy (haichi):

fan works too

OpenStudy (haichi):

D

OpenStudy (rayman5766):

how

OpenStudy (haichi):

C*

OpenStudy (haichi):

soory wrong key

OpenStudy (haichi):

sorry*

OpenStudy (haichi):

-2*8=-16 -16+15=1

OpenStudy (rayman5766):

I just want to know how just in case u try to screw me over also choos any of the that's correct

OpenStudy (haichi):

Oh no im not that kind of person!

OpenStudy (haichi):

It is C thats the correct one for sure

OpenStudy (rayman5766):

ok choose one more also I have to leavee to chat again

OpenStudy (haichi):

k

OpenStudy (haichi):

A and C